# Self-dual gravity in de Sitter space: lightcone ansatz and static-patch scattering

### 07 Jan 2024 v2

@@ II.2.2 Geometry and dynamical data on the past and future horizons
With the above assumptions on the intersection 2-sphere, we can now draw e.g. the future horizon by parallel-transporting the lightlike normal $\ell\equiv\partial_{u}$ along itself, while keeping the $v$ coordinate fixed. The future horizon is then the lightlike hypersurface $v=0$, coordinatized by $(u,\mathbf{\hat{r}})$ with $u\geq 0$, where the lightrays are the lines of constant $\mathbf{\hat{r}}$, and $u$ is an affine coordinate along them. These properties are encoded by the following metric elements at $v=0$:
$g_{\ell\ell}=g_{\ell m}=g_{\ell\bar{m}}=0;\quad g_{\ell n}=\frac{1}{2}\ ,$ (6)
where the constancy of $g_{\ell n}$ encodes the affine nature of the lightlike vector $\ell$ (and its associated coordinate $u$). Similarly, we draw the past horizon at $(u=0,v\geq 0)$, with:
$g_{nn}=g_{nm}=g_{n\bar{m}}=0;\quad g_{\ell n}=\frac{1}{2}$ (7)
@@ V.3 Right-handed initial data on the past horizon, and constraints on the intersection 2-sphere
To obtain the metric components in the $(\ell,m,\bar{m},n)$ basis, we should contract (82) with the basis vectors (80)-(81), and express the result in terms of the initial data $u^{-1}\phi$ (c.f. (70)) and its derivatives $\partial_{u},\partial_{v},\partial_{\mathbf{\hat{r}}}$ along these basis vectors. Using the small-$u$ relation (62) between these derivatives and the Poincare-coordinate gradients $\partial_{\alpha\dot{\alpha}}$, it’s easy to evaluate 8 of the 10 metric components at $u=0$:
$g_{nn}=g_{nm}=g_{n\bar{m}}=g_{\bar{m}\bar{m}}=g_{\ell\bar{m}}=0\ ;\quad g_{\ell n}=g_{m\bar{m}}=\frac{1}{2}\ ;\quad g_{mm}=4\langle\chi q\rangle^{4}\partial_{v}^{2}\frac{\phi}{u}\ ,$ (83)
while $g_{\ell\ell}$ and $g_{\ell m}$ are more complicated. For all the components (including $g_{\ell\ell}$ and $g_{\ell m}$), we notice two features. First, the components are all finite or zero at $u=0$: negative powers of $u$ always come with a contraction of $\bar{\chi}_{\dot{\alpha}}$ with itself, which vanishes. Second, any $\partial_{u}$ derivatives end up multiplied by positive powers of $u$, so they do not appear at $u=0$. Thus, all the components are expressed in terms of the initial data (70) on the past horizon (i.e. at $u=0$), and its derivatives $\partial_{v},\partial_{\mathbf{r}}$ (at most two) along the horizon directions.
From (83), we immediately see that the constraints (7) on the past horizon are satisfied. As we’ll now show, the constraints (3)-(5) on the intersection 2-sphere $u=v=0$ are satisfied as well, assuming the conditions (72) on our initial data. These do not follow directly from (83), because they involve also $g_{\ell\ell}$ and $g_{\ell m}$. Instead, we proceed as follows. First, the constraints are of course satisfied for the pure $dS_{4}$ metric $-2u^{2}\epsilon_{\alpha\beta}\epsilon_{\dot{\alpha}\dot{\beta}}$, as can be checked directly. Now, as discussed in the previous paragraph, the deviation of the metric components from their pure $dS_{4}$ values at $u=0$ are expressed in terms of the initial data $u^{-1}\phi$ and its (at most second) derivatives $\partial_{v},\partial_{\mathbf{r}}$ along the past horizon. But the conditions (72) ensure that at $u=v=0$, the data $u^{-1}\phi$ and its $\partial_{v},\partial_{\mathbf{r}}$ derivatives up to third order all vanish. Therefore, the value and first $\partial_{v}$ derivative of the metric components at $u=v=0$ are the same as in pure $dS_{4}$, thus ensuring the constraints (3),(5).
@@
$\partial_{u}g_{nn}=\partial_{u}g_{nm}=\partial_{u}g_{n\bar{m}}=\partial_{u}g_{\bar{m}\bar{m}}=\partial_{u}g_{\ell\bar{m}}=\partial_{u}g_{\ell n}=\partial_{u}g_{m\bar{m}}=0\ ;$ (84)
@@ V.5 Future horizon data, one lightray at a time
Let us then choose $q^{\alpha}$ normalized as $\hat{t}_{\alpha\dot{\alpha}}q^{\alpha}\bar{q}^{\dot{\alpha}}=-1$, and focus on the particular lightray (with respect to the pure $dS_{4}$ metric) parameterized by (61) with $(\chi^{\alpha},\bar{\chi}^{\dot{\alpha}})=(q^{\alpha},\bar{q}^{\dot{\alpha}})$, i.e. defined by $(v,\mathbf{\hat{r}})=(0,\bm{\sigma}^{\alpha\dot{\alpha}}q_{\alpha}\bar{q}_{\dot{\alpha}})$. We claim that this remains a lightray also in the deformed metric (16), and that $u$ remains an affine parameter along it.
To see that this is true, we begin by writing the basis vectors (79) on our lightray:
$\ell^{\alpha\dot{\alpha}}=\frac{2}{u^{2}}q^{\alpha}\bar{q}^{\dot{\alpha}}\ ;\quad m^{\alpha\dot{\alpha}}=\frac{1}{u}q^{\alpha}q^{\beta}\hat{t}_{\beta}{}^{\dot{\alpha}}\ ;\quad\bar{m}^{\alpha\dot{\alpha}}=\frac{1}{u}\bar{q}^{\dot{\alpha}}\bar{q}^{\dot{\beta}}\hat{t}^{\alpha}{}_{\dot{\beta}}\ ;\quad n^{\alpha\dot{\alpha}}=q^{\alpha}\bar{q}^{\dot{\alpha}}+\frac{1}{2}\hat{t}^{\alpha\dot{\alpha}}\ .$ (98)
Contracting with the metric (16), we immediately find that the metric elements $g_{\ell\ell},g_{\ell m},g_{\ell\bar{m}},g_{\ell n}$ are undeformed from their pure-$dS_{4}$ values, so that the constraints (6) are satisfied; in particular, $\ell^{a}$ remains lightlike. We must be careful, though: the constraints (6) imply affinely parameterized lightrays only when they hold on an entire hypersurface, not just on a single ray. Therefore, we still need to demonstrate that the covariant derivative of $\ell^{a}$ along itself vanishes. Luckily, this is easy to show. We write out the covariant derivative as:
$\ell^{b}\nabla_{b}\ell^{a}=\ell^{b}\partial_{b}\ell^{a}+g^{ad}\ell^{b}\ell^{c}\left(\partial_{b}g_{cd}-\frac{1}{2}\partial_{d}g_{bc}\right)\ .$ (99)
Now, the metric deformation in (16) has both its left-handed spinor indices along $q_{\alpha}$. Since the metric-dependent terms in (99) have at least one of the metric’s indices contracted with $\ell^{\alpha\dot{\alpha}}\sim q^{\alpha}\bar{q}^{\dot{\alpha}}$, we conclude that the covariant derivative (99) remains underformed.
Having established that we’re indeed dealing with an affinely-parameterized lightray of the future horizon, we can proceed to read off the final data on it. Contracting $\bar{m}^{\alpha\dot{\alpha}}$ from (98) with the metric (16), we get:
$g_{\bar{m}\bar{m}}=\frac{1}{4u}(q^{\alpha}q^{\dot{\alpha}}\partial_{\alpha\dot{\alpha}})^{2}\phi-\frac{1}{2}q^{\alpha}q^{\dot{\alpha}}\partial_{\alpha\dot{\alpha}}\phi\ ,$ (100)
where we recall the relation $u=-1/t$. Noticing $q^{\alpha}\bar{q}^{\dot{\alpha}}=(u^{2}/2)\ell^{\alpha\dot{\alpha}}$ and $\ell^{\alpha\dot{\alpha}}\partial_{\alpha\dot{\alpha}}=-2\partial_{u}$, the metric component (100) becomes:
$g_{\bar{m}\bar{m}}=\frac{u^{3}}{4}\partial_{u}^{2}\phi+u^{2}\partial_{u}\phi\ .$ (101)
Taking two more $\partial_{u}$ derivatives, we obtain the final data for the right-handed Weyl curvature component $C_{\ell\bar{m}\ell\bar{m}}$ in terms of the right-handed degree of freedom $\phi(x)$:
$C_{\ell\bar{m}\ell\bar{m}}(u,\mathbf{\hat{r}})=\left(\frac{1}{4}u^{3}\partial_{u}^{4}+\frac{5}{2}u^{2}\partial_{u}^{3}+\frac{11}{2}u\partial_{u}^{2}+2\partial_{u}\right)\phi(x)\ .$ (102)
This relation is morally similar to, though a bit more complicated than, its counterpart (86) on the past horizon.
Finally, we turn to the left-handed final data $c_{\ell m\ell m}(u,\mathbf{\hat{r}})$ in terms of our left-handed degree of freedom $\psi(x)$. As in section V.4, we plug into (23),(48),(52) to translate between $c_{\ell m\ell m}$, $\Psi^{\alpha\beta\gamma\delta}$ and $\psi$, and we again find that only the pure-$dS_{4}$ part of the curvature (52) contributes. Altogether, the left-handed final data reads:

### 10 Mar 2024 v3

@@ II.2.2 Geometry and dynamical data on the past and future horizons
With the above assumptions on the intersection 2-sphere, we can now draw e.g. the future horizon by parallel-transporting the lightlike normal $\ell\equiv\partial_{u}$ along itself, while keeping the $v$ coordinate fixed. The future horizon is then the lightlike hypersurface $v=0$, coordinatized by $(u,\mathbf{\hat{r}})$ with $u\geq 0$, where the lightrays are the lines of constant $\mathbf{\hat{r}}$, and $u$ is an affine coordinate along them. These properties can be encoded by the following relations at $v=0$:
$g_{\ell\ell}=g_{\ell m}=g_{\ell\bar{m}}=0\ ;\quad\partial_{v}g_{\ell\ell}=0\ ;\quad g_{\ell n}=\frac{1}{2}\ ,$ (6)
where the last two conditions ensure the affine nature of the lightlike vector $\ell$ (and its associated coordinate $u$). Similarly, we draw the past horizon at $(u=0,v\geq 0)$, with:
$g_{nn}=g_{nm}=g_{n\bar{m}}=0\ ;\quad\partial_{u}g_{nn}=0\ ;\quad g_{\ell n}=\frac{1}{2}$ (7)
@@ V.3 Right-handed initial data on the past horizon, and constraints on the intersection 2-sphere
To obtain the metric components in the $(\ell,m,\bar{m},n)$ basis, we should contract (82) with the basis vectors (80)-(81), and express the result in terms of the initial data $u^{-1}\phi$ (c.f. (70)) and its derivatives $\partial_{u},\partial_{v},\partial_{\mathbf{\hat{r}}}$ along these basis vectors. Using the small-$u$ relation (62) between these derivatives and the Poincare-coordinate gradients $\partial_{\alpha\dot{\alpha}}$, it’s easy to evaluate 8 of the 10 metric components at $u=0$, along with the $\partial_{u}$ derivative of the component $g_{nn}$:
$g_{nn}=g_{nm}=g_{n\bar{m}}=g_{\bar{m}\bar{m}}=g_{\ell\bar{m}}=0\ ;\quad g_{\ell n}=g_{m\bar{m}}=\frac{1}{2}\ ;\quad g_{mm}=4\langle\chi q\rangle^{4}\partial_{v}^{2}\frac{\phi}{u}\ ;$ (83)
$\partial_{u}g_{nn}=0\ .$ (84)
The remaining components $g_{\ell\ell},g_{\ell m}$ at $u=0$ are more complicated. However, for all the components (including $g_{\ell\ell}$ and $g_{\ell m}$), we notice two features. First, the components are all finite or zero at $u=0$: negative powers of $u$ always come with a contraction of $\bar{\chi}_{\dot{\alpha}}$ with itself, which vanishes. Second, any $\partial_{u}$ derivatives end up multiplied by positive powers of $u$, so they do not appear at $u=0$. Thus, all the components are expressed in terms of the initial data (70) on the past horizon (i.e. at $u=0$), and its derivatives $\partial_{v},\partial_{\mathbf{r}}$ (at most two) along the horizon directions.
From (83)-(84), we immediately see that the constraints (7) on the past horizon are satisfied. As we’ll now show, the constraints (3)-(5) on the intersection 2-sphere $u=v=0$ are satisfied as well, assuming the conditions (72) on our initial data. These do not follow directly from (83), because they involve also $g_{\ell\ell}$ and $g_{\ell m}$. Instead, we proceed as follows. First, the constraints are of course satisfied for the pure $dS_{4}$ metric $-2u^{2}\epsilon_{\alpha\beta}\epsilon_{\dot{\alpha}\dot{\beta}}$, as can be checked directly. Now, as discussed in the previous paragraph, the deviation of the metric components from their pure $dS_{4}$ values at $u=0$ are expressed in terms of the initial data $u^{-1}\phi$ and its (at most second) derivatives $\partial_{v},\partial_{\mathbf{r}}$ along the past horizon. But the conditions (72) ensure that at $u=v=0$, the data $u^{-1}\phi$ and its $\partial_{v},\partial_{\mathbf{r}}$ derivatives up to third order all vanish. Therefore, the value and first $\partial_{v}$ derivative of the metric components at $u=v=0$ are the same as in pure $dS_{4}$, thus ensuring the constraints (3),(5).
@@
$\partial_{u}g_{nm}=\partial_{u}g_{n\bar{m}}=\partial_{u}g_{\bar{m}\bar{m}}=\partial_{u}g_{m\bar{m}}=0\ ;$ (85)
@@ V.5 Future horizon data, one lightray at a time
Let us then choose $q^{\alpha}$ normalized as $\hat{t}_{\alpha\dot{\alpha}}q^{\alpha}\bar{q}^{\dot{\alpha}}=-1$, and focus on the particular lightray (with respect to the pure $dS_{4}$ metric) parameterized by (61) with $(\chi^{\alpha},\bar{\chi}^{\dot{\alpha}})=(q^{\alpha},\bar{q}^{\dot{\alpha}})$, i.e. defined by:
$(v,\mathbf{\hat{r}})=(0,\bm{\sigma}^{\alpha\dot{\alpha}}q_{\alpha}\bar{q}_{\dot{\alpha}})\ .$ (99)
It’s easy to check that this remains a lightray also in the deformed metric (16), and that $u=-1/t$ remains an affine parameter along it. On this lightray, the basis vectors (79) read:
$\ell^{\alpha\dot{\alpha}}=2t^{2}q^{\alpha}\bar{q}^{\dot{\alpha}}\ ;\quad m^{\alpha\dot{\alpha}}=-tq^{\alpha}q^{\beta}\hat{t}_{\beta}{}^{\dot{\alpha}}\ ;\quad\bar{m}^{\alpha\dot{\alpha}}=-t\bar{q}^{\dot{\alpha}}\bar{q}^{\dot{\beta}}\hat{t}^{\alpha}{}_{\dot{\beta}}\ .$ (100)
It is now straightforward to extract the left-handed final data $c_{\ell m\ell m}(u,\mathbf{\hat{r}})$ in terms of our left-handed degree of freedom $\psi(x)$. As in section V.4, we plug into (23),(48),(52) to translate between $c_{\ell m\ell m}$, $\Psi^{\alpha\beta\gamma\delta}$ and $\psi$, and we again find that only the pure-$dS_{4}$ part of the curvature (52) contributes. Altogether, the left-handed final data reads:
@@
With the right-handed final data, we must be more careful. In particular, we can’t directly use the formulas of section II.2: those rely on coordinates adapted to the entire final horizon, whereas the lightcone ansatz is only adapted to the single lightray (99). Of course, we can always just directly compute the Weyl curvature of the metric (16), and extract the appropriate component $C_{\ell\bar{m}\ell\bar{m}}$. An easier path is to note that the lightcone ansatz provides easy access to a different null hypersurface contaning our lightray: the Poincare-coordinate “null hyperplane” $q_{\alpha}\bar{q}_{\dot{\alpha}}x^{\alpha\dot{\alpha}}=0$. We can then extract $C_{\ell\bar{m}\ell\bar{m}}$ from the shear and expansion of this “hyperplane”. This can be further simplified by recalling that the Weyl curvature is conformal. We can rescale the metric (16) to obtain a new metric $\check{g}_{ab}\equiv t^{2}g_{ab}$:
$\check{g}_{\alpha\dot{\alpha}\beta\dot{\beta}}=-2\epsilon_{\alpha\beta}\epsilon_{\dot{\alpha}\dot{\beta}}-q_{\alpha}q_{\beta}q_{\gamma}q_{\delta}\left(t\partial^{\gamma}{}_{\dot{\alpha}}\partial^{\delta}{}_{\dot{\beta}}-2\hat{t}^{\gamma}{}_{(\dot{\alpha}}\partial^{\delta}{}_{\dot{\beta})}\right)\phi(x)\ .$ (102)
With this rescaled metric, the null hyperplane $q_{\alpha}\bar{q}_{\dot{\alpha}}x^{\alpha\dot{\alpha}}=0$ has vanishing expansion and left-handed shear (just like the horizon in the original metric). We can define a vector basis on the hyperplane that plays the same role as the basis $(\ell,m,\bar{m})$ on the horizon:
ˇ$\check{\ell}^{\alpha\dot{\alpha}}=2q^{\alpha}\bar{q}^{\dot{\alpha}}\ ;\quad\check{m}^{\alpha\dot{\alpha}}=q^{\alpha}q^{\beta}\hat{t}_{\beta}{}^{\dot{\alpha}}\ ;\quad\check{\bar{m}}^{\alpha\dot{\alpha}}=\bar{q}^{\dot{\alpha}}\bar{q}^{\dot{\beta}}\hat{t}^{\alpha}{}_{\dot{\beta}}\ .$ (103)
Specifically, these basis vectors Lie-commute, $\check{\ell}$ is an affine null generator with respect to the metric (102), $\check{m}$ is a left-handed null vector, and the inner product of $\check{m},\check{\bar{m}}$ under the metric (102) is $\check{g}_{m\bar{m}}=\frac{1}{2}$. We can now extract the Weyl component $\check{C}_{\ell\bar{m}\ell\bar{m}}$ for the metric (102) as:
$\check{g}_{\bar{m}\bar{m}}$ $=g_{\bar{m}\bar{m}}=-\frac{1}{4}\left(t(q^{\alpha}\bar{q}^{\dot{\alpha}}\partial_{\alpha\dot{\alpha}})^{2}+2q^{\alpha}\bar{q}^{\dot{\alpha}}\partial_{\alpha\dot{\alpha}}\right)\phi(x)\ ;$ (104)
$\check{C}_{\ell\bar{m}\ell\bar{m}}$ =(ˇ$=(\check{\ell}^{a}\partial_{a})^{2}\check{g}_{\bar{m}\bar{m}}=-\frac{t}{4}(q^{\alpha}\bar{q}^{\dot{\alpha}}\partial_{\alpha\dot{\alpha}})^{4}\phi(x)\ .$ (105)
Rescaling back into the metric (16) and basis (100), we finally obtain the right-handed data on our chosen lightray of the future horizon as:
$C_{\ell\bar{m}\ell\bar{m}}=t^{4}\check{C}_{\ell\bar{m}\ell\bar{m}}=\frac{(q^{\alpha}\bar{q}^{\dot{\alpha}}\partial_{\alpha\dot{\alpha}})^{4}\phi(x)}{4u^{5}}\ .$ (106)
We note the simplicity of this formula (as compared e.g. with the metric component (104)), as well as its similarity to the initial right-handed data (87) and the final left-handed data (101).